OpenStudy (anonymous):

A candle is 10 in. tall after burning for 2 hours. After 3 hours, it is 8.5 in. tall. How tall will the candle be after 6 hours? 22 in. 2 in. 1.5 in. 4 in.

OpenStudy (anonymous):

4 inches.

OpenStudy (anonymous):

What a tricky question. I like how they try to stump people here but in my eyes its a great disappointment

OpenStudy (anonymous):

Oops, the points were supposed to be (2, 10) and (3, 8.5)

OpenStudy (anonymous):

8.5 - 10/3 - 2 -1.5/1 -1.5 This is your slope.

OpenStudy (anonymous):

y = -1.5x + 13 This equation goes through both points (2, 10) and (3, 8.5)

OpenStudy (anonymous):

Plug in 6 into your equation and solve for y: y = -1.5x + 13 y = -1.5(6) + 13 y = -9 + 13 y = 4 So we have (6, 4) which means the candle will be 4 inches tall.

OpenStudy (anonymous):

I support this answer^